Calculate the MP for a lease (annuity due) for $125,500 with terms 6%, 5 years.

(a) Payments are made at the beginning of each month and the Expected Salvage Value is zero

(b) Salvage Value is now $25,000
Answer #1

HI,

Please see below:

Answer 1:

$125,500.00 Loan Amount Number of Loan Years APR Monthly Rate Monthly Payments 6.00% 0.50% $2,414.20

125500 Loan Amount Number of Loan Years APR Monthly Rate Monthly Payments 0.06 =C4/12 --PMT(C5,C3*12,C2,,1)

Answer 2:

$106,818.55 Salvage value after 5 years $25,000.00 PV of Salvage $18,681.45 Loan Amount Number of Loan Years APR Monthly Rate

Salvage value after 5 years 25000 PV of Salvage =12/(1+F4)^F3 Loan Amount T=125500-13 Number of Loan Years APR 0.06 Monthly R